OpenStudy (anonymous):

what is 15/50 in lowest terms? A. 1/2 B. 3/10 C. 1/4 D. 3/8

OpenStudy (anonymous):

\[ \frac{15}{50} = \frac{3 \times 5}{10 \times 5} = \frac{3}{10} \]

OpenStudy (anonymous):

i got it right the first time

OpenStudy (anonymous):

Well gj then =]